# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

### v2

@@ A.2 Convex-set valued functions in the complex setting
Lemma 3.13 in [Cruz_Uribe_Bownik_Extrapolation] yields, for a given measurable real vector valued function $f,$ a measurable real convex body valued function $F$ with $|F(x)|=|f(x)|$ at every $x\in\Omega.$ For the complex version, we consider a measurable complex vector valued function $f$ and we construct the corresponding measurable complex convex body valued function $F$ with $|F(x)|=|f(x)|$ at every point in the domain. This is achieved performing one major change that is a recurring theme in the passage from the real to the complex case. Namely, given $f\in\mathrm{L}^{1}(\Omega,\mathbb{C}^{d})$, one defines the convex-set valued map $F$ by
$F(x):=\{zf(x):\leavevmode\nobreak\ z\in\mathbb{C}\text{ with }|z|\leq 1\},\quad x\in\Omega.$
@@
This map is measurable just as in the real case in [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.13] because also the closed unit disk in the complex plane has a dense countable subset.
The validity of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.14], which states that the Aumann integral of a closed set valued function with respect to a nonatomic measure is a convex (not necessarily closed) set, is obvious in the complex case, because this theorem concerns only the topological and real vector space $\mathbb{R}^{d}$, remaining valid for $\mathbb{R}^{2d}$ and thus also $\mathbb{C}^{d}$.
To obtain a complex version of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.15], which states that the closure of the Aumann integral of a convex body valued function is also a convex body, we need once again to directly replace $\mathbb{R}^{d}$ by $\mathbb{C}^{d}$. The notion of real symmetric sets is accordingly replaced by the notion of complex symmetric sets. The Dunford–Pettis theorem and Mazur’s lemma hold equally well with the complex numbers as the underlying field of scalars.
The statement and proof of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.16] remain true verbatim in the complex case with $\mathbb{R}^{d}$ replaced by $\mathbb{C}^{d}$, because as mentioned Mazur’s lemma remains true in complex Banach spaces.
It is obvious that [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.17, Corollary 3.18], which deal with linearity and monotonicity of Aumann integrals of convex body valued functions, and their proofs remain both true verbatim in the complex case with $\mathbb{R}^{d}$ replaced by $\mathbb{C}^{d}$.
Observe again that the definition of the Aumann integral makes use only of the topological space and real vector space structures of $\mathbb{R}^{d}$. Note also that complex symmetric sets are in particular real symmetric sets. Since [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.19], which states that the Aumann integral of a convex body valued function $F$ equals $\{0\}$ only if $F(x)=0$ almost everywhere, holds equally well for $\mathbb{R}^{2d}$, we deduce that it remains true also for $\mathbb{C}^{d}$.
Finally, in the complex versions of [Cruz_Uribe_Bownik_Extrapolation, Proposition 3.20, Proposition 3.21], which are respectively analogues of Hölder’s inequality and Minkowski’s inequality for Aumann integrals of convex body valued functions, one replaces $\mathbb{R}^{d}$ by $\mathbb{C}^{d}$ and then considers complex-symmetric sets instead of real-symmetric ones and complex homogeneous norms on $\mathbb{C}^{d}$ instead of real homogeneous norms on $\mathbb{R}^{d}$. The statements and proofs remain in both cases otherwise completely unchanged, because complex symmetric sets are in particular real symmetric sets, complex homogeneous norms are in particular real homogeneous, and as remarked above, [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.17] extends obviously to the complex case.
@@ A.3 Seminorm functions
In this last section, we cover the adaptation of [Cruz_Uribe_Bownik_Extrapolation, Section 4] to the complex setting. As mentioned in Subsection A.1, the definition of seminorm functions itself only requires substituting the vector space $\mathbb{R}^{d}$ by $\mathbb{C}^{d}$ and real homogeneity by complex homogeneity.
For the statement and proof of [Cruz_Uribe_Bownik_Extrapolation, Theorem 4.2], that gives a one-to-one correspondence between seminorm functions and convex body valued functions, one only needs the obvious modifications. This is, one needs to substitute $\mathbb{R}^{d}$ by $\mathbb{C}^{d},$ consider the Borel $\sigma$-algebra $\mathcal{B}$ on $\mathbb{C}^{d},$ use countable dense sets of $\mathbb{C}^{d}$ such as $\mathbb{Q}(i)^{d}$ and use the complex version of Hahn-Banach Theorem.
Regarding [Cruz_Uribe_Bownik_Extrapolation, Lemma 4.4], which gives that every measurable convex body valued mapping is the pointwise limit of simple measurable mappings (in the Hausdorff topology), both its statement and its proof follow verbatim under the substitution of $\mathbb{R}^{d}$ by $\mathbb{C}^{d}.$ Note that the characterization of the convergence of convex bodies used in [Cruz_Uribe_Bownik_Extrapolation] can be used in the same way in our context because it actually applies to general nonempty compact convex sets (see [Schneider_Convex_Bodie_1993, Theorem 1.8.7]).
To get [Cruz_Uribe_Bownik_Extrapolation, Theorem 4.5], which characterises integrably boundedness of a measurable convex body valued function in terms of the corresponding seminorm function, one needs to follow the same argument replacing $\mathbb{R}^{d}$ by $\mathbb{C}^{d}$ and keeping in mind that we have changed the definition of function $h_{F(x)}(v)$ to
$h_{F(x)}(v)\coloneq\sup_{w\in F(x)}|\langle v,w\rangle|$
Next, all results about $\mathrm{L}^{p}$ spaces of convex-set valued functions follow with the replacement of $\mathbb{R}^{d}$ by $\mathbb{C}^{d},$ since these are based on the topology given by the Hausdorff metric.
Finally, all considerations about matrix weights and seminorms also hold. Here, in addition to the substitution of $\mathbb{R}^{d}$ by $\mathbb{C}^{d},$ it is also necessary to replace orthogonal matrices by unitary ones and symmetric matrices by those that are hermitian.
@@ A.4 Main results including the Extrapolation Theorem with matrix weights
In the previous subsections we have explained how the theory developed to deal with real valued matrix weights in [Cruz_Uribe_Bownik_Extrapolation, Sections 2–4] can be modified to be applied to complex valued matrix weights. Once these tools have been conveniently adapted, they can be used to get the complex version of the main results in that article without further changes. In other words, the exposition of [Cruz_Uribe_Bownik_Extrapolation, Sections 5–9] holds for the complex setting by using the results explained in the current appendix. In particular, Theorem A holds.
[…]
Spyridon Kakaroumpas
Würzburg Mathematics Center of Communication and Interaction,
Julius-Maximilians-Universität Würzburg,
Campus Hubland Nord,
Emil-Fischer-Straße 41,
97074 Würzburg, Germany
E-mail address:
spyridon.kakaroumpas@uni-wuerzburg.de
Odí Soler i Gibert
Universitat Politècnica de Catalunya - BarcelonaTech (UPC),
Pavelló I,
Diagonal 647,
08028 Barcelona, Catalunya
E-mail address:
odi.soler@upc.edu

### v3

@@ A.2 Convex-set valued functions in the complex setting
Finally, another main point in [5, Section 3] is that Lemma 3.13 in [5] yields, for a given measurable real vector valued function $f,$ a measurable real convex body valued function $F$ with $|F(x)|=|f(x)|$ at every $x\in\Omega.$ For the complex version, we consider a measurable complex vector valued function $f$ and we construct the corresponding measurable complex convex body valued function $F$ with $|F(x)|=|f(x)|$ at every point in the domain. This is achieved performing one major change that is a recurring theme in the passage from the real to the complex case. Namely, given $f\in\mathrm{L}^{1}(\Omega,\mathbb{C}^{d})$, one defines the convex-set valued map $F$ by
$F(x):=\{zf(x):\penalty 10000\ z\in\mathbb{C}\text{ with }|z|\leq 1\},\quad x\in\Omega.$
@@
This map is measurable just as in the real case in [5, Lemma 3.13] because also the closed unit disk in the complex plane has a dense countable subset.
Finally, the rest of the results in [5, Section 3] follow by the general principles mentioned before. Namely, one uses that $\mathbb{C}^{d}$ shares an equivalent structure to $\mathbb{R}^{2d}.$ In addition, all abstract theorems used to treat Aumann integrals are equally valid in the complex setting as in the real one.
@@ A.3 Seminorm functions
The adaptation of [5, Section 4] is performed similarly to that for [5, Section 2] outlined above, using the corresponding results from the previously described adaptation of [5, Section 3]. Additional ingredients are countable dense sets of $\mathbb{C}^{d}$ such as $\mathbb{Q}(i)^{d}$ and the complex version of Hahn-Banach Theorem. Finally, note that the characterization of the convergence of convex bodies used in [5, Section 4] can be used in the same way in our context because it actually applies to general nonempty compact convex sets (see [39, Theorem 1.8.7]).
@@ A.4 Main results including the Extrapolation Theorem with matrix weights
In the previous subsections we have explained how the theory developed to deal with real valued matrix weights in [5, Sections 2–4] can be modified to be applied to complex valued matrix weights. Once these tools have been conveniently adapted, they can be used to get the complex version of the main results in that article without further changes. In other words, the exposition of [5, Sections 5–9] holds for the complex setting by using the results explained in the current appendix. In particular, Theorem A holds.